### v1

# Using $k$-medoids for distributed approximate similarity search with arbitrary distances

Abstract
This paper presents GMASK, a general algorithm for distributed approximate similarity search that accepts any arbitrary distance function. GMASK requires a clustering algorithm that induces Voronoi regions in a dataset and returns a representative element for each region. Then, it creates a multilevel indexing structure suitable for large datasets with high dimensionality and sparsity, usually stored in distributed systems. Many similarity search algorithms rely on $k$-means, typically associated with the Euclidean distance, which is inappropriate for specific problems. Instead, in this work we implement GMASK using $k$-medoids to make it compatible with any distance and a wider range of problems. Experimental results verify the applicability of this method with real datasets, improving the performance of alternative algorithms for approximate similarity search. In addition, results confirm existing intuitions regarding the advantages of using certain instances of the Minkowski distance in high-dimensional datasets.
Keywords Approximate similarity search  $\cdot$ k-medoids  $\cdot$ arbitrary distance function
1 Introduction
The goal of similarity search is to find elements from a dataset, represented by feature vectors in a multidimensional space, similar to a given query object according to a particular similarity or dissimilarity measure. Numerous data indexing methods, collectively known as access methods, have been proposed to speed up this task. Nevertheless, the results remain strongly conditioned by the problems inherent to the representation of datasets in high-dimensionality and high-sparsity spaces. These algorithms are frequently conceived to work with data having spatial components or represented in a vector space with few dimensions. However, they become unusable in high-dimensional spaces, where many features characterise each element due to the Curse of Dimensionality [1]. As a result, the distance between elements tends to become increasingly uniform, making it challenging to differentiate between them based solely on their distances or when the number of dimensions significantly exceeds the quantity of data points. Moreover, several studies indicate that in datasets with over 20 dimensions, exact similarity search methods rarely outperform exhaustive scanning.
Owing to this constraint, novel approaches have emerged in recent years to swiftly identify enough related items within a reasonable time frame compared to linear scanning, often referred to as approximate similarity search. One of the newest approaches is MASK (Multilevel Approximate Similarity search with k-means) [2], a novel indexing method that involves an unconventional application of the $k$-means partitioning algorithm [3, 4] to create a multilevel index structure for approximate similarity search.
Traditionally, the input to $k$-means is always a fixed number of $k$ prototype points or centroids. Then, following an iterative process, the dataset is divided into $k$ groups (Voronoi cells), assigning each point to its closest centroid according to a given distance function, so that each centroid represents all of the elements of that cluster. In contrast, leveraging current powerful computing capabilities, $k$-means can be forced to generate a high number of centroids that inundate the dataset. Then, these centroids become a new dataset to be grouped at the upper level, repeating this process until the complete construction of the search tree is achieved. This bottom-up creation of the multilevel index enables the distribution of this task among different computing nodes, which also store distinct data partitions.
Although it employs an innovative approach that capitalises on the benefits offered by contemporary computing systems and distributed platforms, the performance of this indexing method remains significantly impacted by the inherent constraints of the $k$-means algorithm. The choice of the distance function significantly impacts the outcome of $k$-means clustering, as it determines how the similarity between data points is calculated and can lead to different outcomes in terms of the shape and separation of the clusters.
Until now, experiments conducted with $k$-means have relied on the Euclidean distance to calculate relative distances between points, following the conventional inclination to associate the concept of distance with that used in geospatial contexts. Nonetheless, relying on the Euclidean distance as the default option in indexing structures and algorithms when working with high-dimensional data, despite its widespread use, can lead to sub-optimal results. This is because, in high-dimensional spaces, the discriminating power of the $L_{p}$ norm decreases for higher values of $p$ due to their specific features (such as dispersion, outliers, and concentration of points around certain values), leading to poor differentiation between proximate and distant elements.
This paper introduces GMASK (General Multilevel Approximate Similarity Search with $k$-medoids), a novel algorithm that facilitates the incorporation of arbitrary distances by employing data partitioning algorithms that induce Voronoi regions in a dataset and yield a representative element, exemplified by $k$-medoids. The simple introduction of $k$-medoids turns GMASK into a versatile algorithm capable of adapting to the characteristics of any dataset (high dispersion, high dimensionality, outliers, etc.) and seek for the optimal results. This approach allows using multiple $\ell_{p}$ norms or even other distance functions, making it compatible with a broader spectrum of problems. To comprehensively evaluate its performance, an extensive series of experiments will be conducted across datasets varying in type, size, dimensionality, and sparsity.
This paper adheres to the following structure. Section “Background and related work” reviews previous research related to similarity search. Section “Methodology” describes GMASK, emphasising the potential for enhancing the algorithm’s performance by exploring the application of other distance functions beyond the Euclidean one, like other Minkowski distances. The findings of four empirical experiments carried out over real-world datasets with quite diverse attributes are presented in Section “Experimental results”. These tests compare GMASK performance with other well-known algorithms, such as FLANN and PyNNdescent. Finally, Section “Conclusions and future work” recaps the main conclusions of this work and suggests promising lines for future work.
1.1 Contributions
The potential for achieving improved results, coupled with the lack of prior empirical studies that validate this proposal and the need for learning paradigms applicable to a wider range of dissimilarity measures [5], justifies our efforts to explore alternative Minkowski distances for data indexing. This endeavour is aimed at challenging the misconception we have identified and fostering a deeper understanding of the significant impact of using alternative distance metrics when working with high-dimensional data.
  • GMASK allows the use of different clustering algorithms for the bottom-top construction of the index tree, such as $k$-medoids, a broader form of the $k$-means algorithm that offers parameterisation with several distance metrics, in particular Minkowski distances.
  • Evaluating the performance of different $L_{p}$ norms or Minkowski distances for approximate $k$-nearest neighbour search. This assessment involves testing GMASK against other state-of-the-art methods to illustrate that the selection of a suitable distance function, while considering the features of the target dataset, has a significant impact on performance when searching for reasonably similar regions.
2 Background and related work
This section provides minimal background on the concepts and strategies underlying similarity search methods, including the description of current similarity search algorithms.
2.1 Similarity search
Similarity search, often referred to as proximity search [6, 7, 8], is a cornerstone for applications in many different fields, such as databases [9], information retrieval [10], computer vision [11] and bioinformatics [12, 13], among others. It aims to find elements that are similar (close) to a given query object, according to some measure of similarity or, conversely, dissimilarity.
Thus, given a query object $q\in\mathbb{X}$, the similarity search tackles some of the following scenarios, represented in Figure 1 [7]:
  • Point query: Finding elements within $\mathbb{X}$ that have exactly the same feature values as $q$.
  • Range query: Identifying a subset of elements $\{o_{i}\}\in\mathbb{X}$ whose feature values do not exceed a given similarity threshold $r$ around $q$ such that $\forall\{o_{i}\},\delta(o_{i},q)\leq r$.
  • Nearest-neighbour query: Finding elements in $\mathbb{X}$ whose features are the closest to the feature vector representing $q$. In this case, the goal might be to retrieve the single element (set of $k$ elements) closest to $q$, known as the nearest neighbour ($k$-nearest neighbours) and denoted by $NN_{q}$ ($k-NN_{q}$). Identifying the $k$-nearest neighbours of a given object is one of the most widely used classification methods in machine learning.
Figure 1: Conceptual representation of several types of query searches.
Many techniques for data indexing, collectively known as access methods, can speed up these queries. These methods create an index structure that partitions the domain $\mathbb{X}$ into different regions based on a distance function. Then, we can leverage these regions to solve similarity queries. There is a wide range of access methods, including exact approaches, approximate solutions and variants tailored for distributed computing systems.
2.2 Algorithms for similarity search
Exact similarity search methods identify the completely accurate result set for a query by using the absolute positions of objects to group elements and search for similar instances. These methods are categorised as Point Access Methods (PAM) for point-based elements and Spatial Access Methods (SAM) for elements with spatial dimensions, such as lines, polygons or higher-dimensional polyhedra. The general acronym SAM is commonly used in the research literature to refer to both categories of indexing strategies in vector spaces [10]. These approaches are most effective with elements represented in a relatively small number of dimensions, making Curse of Dimensionality one of the biggest challenges in data science today.
Metric Access Methods (MAM) provide a general framework for similarity search problems within metric spaces [6, 8], exploiting the triangle inequality to partition the metric space into subspaces, effectively filtering out portions of the dataset that cannot contain valid results. On the one hand, MAM exact search can be expensive in terms of computational time and updating the index structure due to dynamic data. On the other hand, MAM indexes have serious problems in efficiently reducing the search for candidate elements in a query when dealing with high-dimensional data. There are many MAM for queries in metric spaces, some of which use clustering algorithms, such as $k$-means [4] or $k$-medoids [14], to recursively partition the entire set of elements creating a hierarchical structure (tree) of nested clusters.
Approximate indexing methods are solutions that strike a balance between accuracy and speed, making them suitable for both vector and metric spaces [15]. Although these methods may introduce some errors into the results, they also provide a faster response time by calculating fewer distances. As a result, these methods prove feasible in scenarios where exact methods cannot provide a fast response, such as in high-dimensional problems or large datasets, where exact methods default to a full scan. Some of the most widely used exact algorithms are compared and summarised in [16], which also introduces a comprehensive taxonomy to classify them according to relevant characteristics.
Rather than computing the exact result set for a query, approximate similarity search algorithms adopt various strategies (e.g., modifying the representation space, reducing the number of pair distances to compute) to identify a reduced set of candidate items that are likely to be the most similar. FastMap, introduced in [17], is one of the first algorithms to be developed for approximate similarity search in metric spaces. It is based on a method to obtain a fast projection of elements in the original dataset into a user-defined $k$-dimensional space, preserving the distances between pairs of elements in order to achieve an approximate response through dimensionality reduction. In this category of algorithms, MASK [2] emerged as a novel indexing approach characterised by a multilevel design tailored for approximate similarity search. It features an innovative utilisation of the $k$-means partitioning algorithm to quickly reduce the number of checked regions when searching for results.
2.3 Using different distance functions
The dissimilarity value captures the notion of closeness between two objects, which is interpreted as a distance in a suitable space. A distance function that avoids introducing any correlation between features, such as the Minkowski distance [18] or $\ell_{p}$ norm, can quantify this value [17]. The Manhattan or city block distance ($\ell_{1}$ norm), the Euclidean distance ($\ell_{2}$ norm), and the Chebyshev distance ($\ell_{\infty}$ norm) are typical instances of this family. As many multidimensional indexing structures were initially proposed in the context of spatial systems, many indexing structures and algorithms have commonly relied on the Euclidean distance, derived from its conventional use in two- or three-dimensional spatial applications.
However, several negative consequences and challenges arise when dealing with high dimensional problems, an effect widely known as the Curse of Dimensionality. As the number of dimensions increases, the volume of the space grows exponentially, and the density of points in the space decreases. This sparsity poses a challenge in accurately measuring distances and similarities between data points, since proximity becomes less meaningful. Consequently, points can be very far apart, even though they are close according to a metric such as Euclidean distance. This effect increases the computational complexity of algorithms and operations performed in high-dimensional spaces, often resulting in slower and less efficient computations. For this reason, several authors have pointed out that, due to the Curse of Dimensionality, it is very costly to find the exact nearest neighbour in high-dimensional Euclidean spaces [19, 20].
As a result, although Euclidean spaces reflect our natural conception of space, this natural assumption is challenging to extrapolate when dealing with high-dimensional data. Moreover, traditional clustering algorithms, such as $k$-means, often specify the Euclidean distance as the only appropriate metric without considering other possibilities. This choice brings in a self-imposed limitation that may produce results far from the best [21, 22, 5].
In recent years, several surveys [23, 24, 25] have addressed the evaluation of various metrics for similarity searches across spaces of diverse topology, a burgeoning topic offering insights into the most suitable similarity measures for different problems and contributing to the ongoing development of new methods for similarity search.
Therefore, we aim to introduce GMASK, an algorithm that, being based on a generalisation of earlier concepts previously introduced by MASK, signifies a substantial departure from the latter. GMASK enables the integration of arbitrary distances through the use of data partitioning algorithms that generate Voronoi regions and assign a prototype for each of these regions. Specifically, the utilisation of the $k$-medoids algorithm allows for the incorporation of multiple $\ell_{p}$ norms, facilitating parametrisation with different distance metrics and the execution of validation experiments. Our findings demonstrate the feasibility of employing diverse distances beyond the Euclidean metric, highlighting the potential for achieving superior results in data indexing, contingent upon the specific characteristics of the dataset.
3 Methodology
This section describes the methodology developed for data indexation using arbitrary distance functions. First, it illustrates the GMASK algorithm as a general approximate indexing method accepting any distance function. Hereafter, it delves into the characteristics of Minkowski distances, emphasising the inherent limitations in $k$-means that justify the implementation of GMASK. Consequently, we justify adopting $k$-medoids, an alternative clustering method that leverages the opportunities provided by Minkowski distances to overcome previous restrictions derived from using the Euclidean distance. The design of experimental tests is described in Subsections “Performance evaluation”, “Benchmark algorithms for similarity search”, and “Datasets”, including metrics to assess performance, state-of-the-art algorithms for comparison, and the datasets selected to gauge effectiveness.
3.1 GMASK
As mentioned above, there is a close connection between machine learning and similarity search indexing. Firstly, the foundations for designing data access methods rely on familiar concepts in machine learning classification, such as dissimilarity spaces and distance functions. Secondly, clustering methods like $k$-means or $k$-medoids are used in many of these algorithms to create indexing structures for similarity search and improve search performance. Besides, they are also affected by the same limitations, such as the Curse of Dimensionality and the challenge of identifying the actual intrinsic dimensionality governing some problems to make them solvable.
The typical choice in partitioning clustering methods is to find a low value of $k$ that generates different groups in the dataset. GMASK follows an unconventional approach for clustering, generating a high number of $k$ representative points for each level. The only requirement in GMASK is using a clustering method that generates Voronoi regions and returns a representative point for each of them. In this implementation, GMASK uses the $k$-medoids clustering method, inducing a Voronoi tessellation and returning an element representing each region. Thus, generating a large number of representative points, GMASK summarises the underlying data points with finer detail. In the same way, this indexing method applies to any dissimilarity space, including the particular case of metric spaces or vector spaces. Thanks to the $k$-medoids adaptation in GMASK, this algorithm presents an important improvement over MASK: it is able to work with different distance functions. This modification brings GMASK the capability of adaptation to the specific characteristics of any dataset. In the “Experimental results” section, we show that this unusual application of clustering has attractive properties and yields superior performance, even with high-dimensional datasets residing in feature spaces that are typically sparse.
At this point, it is essential to note that the goal of GMASK is not to cluster data but to use the partitioning clustering method to construct a multilevel index to find elements. With more powerful computational infrastructures and large memory and storage capacity becoming prevalent, the advantage of this method lies in that previous restrictions on the number of clusters at each level become less relevant. In this new scenario, we can exploit the effect of introducing a high number of prototype points into the dataset, leveraging the capabilities of the computing infrastructure.
Figure 2 illustrates the design principles of GMASK. At the lowest level, we have the data points in $\mathcal{X}$ to be indexed. The blue boxes in the diagram represent different data partitions or groups that could be stored on various computing nodes. The novelty lies in creating the multilevel structure of $k$ medoids, following a bottom-up approach. The first level of medoids (identified as $n_{1}$ in Figure 2) summarises the data points at the lowest level so that points assigned to each cluster are represented by its prototype (medoid). The second level (denoted $n_{2}$) corresponds to medoids representing the prototypes computed at the first level, $n_{1}$. The same strategy is repeated, creating additional prototype layers that summarise the immediately lower group of medoids. The procedure stops when the number of medoids at the top level of this index structure becomes manageable.
Figure 2: Conceptual representation of the updated multilevel structure approach created by GMASK.
The process of constructing this multilevel index structure depends on two initial parameters:
  • The group length quantifies the number of elements in each data partition at the lowest level, configured according to the available computing power (and the total number of nodes in distributed computing systems). For the baseline, we assume that all partitions are of the same size, although this condition is not strictly necessary.
  • The number of medoids calculated for each group. At this point, the question of selecting the most optimal number of medoids has not yet been addressed, but it is not very important. Contrary to the standard procedure, our approach suggests that, for information indexing, one should use as many medoids as the computational infrastructure can afford.
The clustering at the first level reduces the initial number of data points by a particular proportion, the data summarisation ratio, which is determined by the ratio between the group length and the number of medoids. This ratio also determines the number of layers of the multilevel index and the stop condition when the number of medoids at the top layer becomes lower than the group length.
Each medoid is a prototypical point representing all the elements assigned to its cluster. The number of points assigned to each cluster may vary depending on their distance from the corresponding medoid. For this reason, regions with higher data density will also tend to receive more medoids. By generating a high number of medoids, we can capture finer details of the underlying data, allowing us to index not only high-density regions but also outliers and clusters of elements distinct from the main concentrations of points. Algorithm 1 describes the construction of this multilevel index structure.
Algorithm 1 Multilevel index construction in GMASK
1:function Clustering($distFunc,nPrototypes,points$)
2:     $\triangleright$ A clustering method returning Voronoi regions
3:     $\triangleright$ and a prototype point representing each region
4:end function
5:
6:procedure GMASK($data,lengthGroup,nPrototypes$)
7:     $ngroups\leftarrow lengthData/lengthGroup$;
8:     $vector\leftarrow$ split($data,nGroups$); $\triangleright$ Divide dataset into nGroups
9:     $\triangleright$ Initialise lists $layerPoints$, $layerLabels$, and variable $idLayer$
10:     while $ngroups\geq 1$ do
11:         $\triangleright$ Initialise groupsPoints, groupsLabels and points
12:         for $idGroup=1,nGroups$ do
13:              $points\leftarrow vector[idGroup]$;
14:              $groupsPoints.add($clustering($distFunc,nPrototypes,points$));
15:              $groupsLabels.add(clusters.labels)$;
16:         end for
17:         $layerPoints.add(groupsPoints)$;
18:         $layerLabels.add(groupsLabels)$;
19:         $ngroups\leftarrow length(groupsPoints)/lengthGroup$;
20:         if $ngroups\geq 1$ then
21:              $vector\leftarrow$ split($layerPoints,nGroups$);
22:         end if
23:         $idLayer\mathrel{+}=1$;
24:     end while
25:     $layers\leftarrow idLayer-1$;
26:end procedure
Once built according to the rules described above, the index is ready to accept point queries, which systematically search for all points in the original dataset, seeking a perfect match. To solve any query using this multilevel index, we start at the top level of the medoids hierarchy, steering the search according to the distance between the medoids at each level and the target query object $q\in\mathcal{X}$.
3.2 Dissimilarity measures
A dissimilarity space [26] defines the rules for measuring the distance between elements represented in a feature space. As discussed in Section “Similarity search”, similarity captures the notion of closeness between two objects. Euclidean vector spaces are widely applied in previous research, as they reflect and intuitive interpretation of a space: vectors are elements that can be added, multiplied by a scalar or projected onto each other.
Being $\mathbb{X}$ a set of elements, the function
$\delta:\mathbb{X}\times\mathbb{X}\to\mathbb{R}$ (1)
returns a real number $\delta(a,b)$ that shall be used to quantify the difference between two or more elements contained on this set, denoting a measure of distance.
Distance functions have the following properties:
  • $(p1)\quad\forall a,b\in\mathbb{X},\delta(a,b)\geq 0$ positiveness,
  • $(p2)\quad\forall a,b\in\mathbb{X},\delta(a,b)=\delta(b,a)$ symmetry,
  • $(p3)\quad\forall a\in\mathbb{X},\delta(a,a)=0$ reflexivity,
  • and in most cases
  • $(p4)\quad\forall a,b\in\mathbb{X},a\neq b\to\delta(a,b)>0$ strict positiveness.
If it also satisfies
  • $(p5)\quad\forall a,b,c\in\mathbb{X},\delta(a,b)\leq\delta(a,c)+\delta(c,b)$, commonly known as triangle inequality,
$\delta$ is indeed a metric, and the pair $(\mathbb{X},\delta)$ is called a metric space [6]. If the distance does not satisfy the strict positiveness property $(p4)$ then the space is called a pseudo-metric space, as well as if it does not hold the symmetry property $(p2)$, we may have a quasi-metric space. Nevertheless, there are some primitive spaces where these metric constraints are not necessarily satisfied. We also use the word ’dissimilarity’ when referring to this difference or distance between elements [18].
Normed and inner product spaces are special cases of metric spaces, where a metric is defined either by a norm or an inner product. An Euclidean space is a specific case of a Hilbert space. In turn, a Hilbert space is an inner product space and a particular case of a Banach space, an example of normed spaces in which a norm metric $\ell norm$ has been defined [5].
3.3 Minkowski distance
Minkowski distances or $\ell_{p}norms$ are distance measures commonly used to compute the dissimilarity of objects described by numerical attributes. However, dissimilarity can also be quantified using non-numerical attributes through other alternatives, such as the edit distance widely used for text attributes.
Let $\ell_{\infty}^{p},p\geq 1$ be a vector space of real sequences such that $\sum_{i=1}^{\infty}|x_{i}|^{p}$ with the norm given by $||x||_{p}(\sum_{i=1}^{\infty}|x_{i}|^{p})^{\frac{1}{p}}$. This norm induces the $\ell norm$ or Minkowski metric $d_{p}$ and consequently $\ell_{p}^{m}$ and $\ell_{p}^{m}=(\mathbb{R}^{m},d_{p})$ are Banach spaces.
Thus, assuming $x,y\in\mathbb{R}^{m}$, the $\ell norm$ or Minkowski distance is defined as follows:
$d_{p}(x,y)=(\sum_{i=1}^{m}|x_{i}-y_{i}|^{p})^{\frac{1}{p}},p\geq 1$ (2)
where depending on the $p$ value we can induce:
  • $d_{1}(x,y)=\sum_{i=1}^{m}|x_{i}-y_{i}|$, the Manhattan or city-block distance.
  • $d_{2}(x,y)=d_{E}(x,y)=(\sum_{i=1}^{m}|x_{i}-y_{i}|^{2})^{\frac{1}{2}}$, the Euclidean distance.
[…]
Minkowski distances with $p<1$ cannot be considered as proper metrics since they do not hold the triangle inequality. Nonetheless, several authors have shown that fractional distances can significantly improve the effectiveness of standard clustering algorithms such as $k$-means [22].
3.4 Limitations of k-means and the Euclidean distance
In unsupervised machine learning with $k$-means, the user must determine the appropriate number of clusters for a given problem, and a centroid represents each cluster. Carefully choosing the number of centroids is crucial because too many centroids would not effectively summarise the underlying data, and too few centroids would result in grouping unrelated data points. However, the locations provided to these initial centroids can greatly impact the final outcome, making it a sensitive aspect.
To address this limitation, algorithms such as $k$-means++ [27] propose an approach where the initial random centroids are spread more evenly, leading to improved results. Nevertheless, although several more powerful $k$-means implementations have tried to reduce the performance impact resulting from sub-optimal initial centroid placement, only a few have explored the potential for achieving better solutions by choosing a more appropriate distance.
Due to the nature of the $k$-means algorithm, the choice of distance function can also exert a significant impact on $k$-means outcomes, as it determines how similarity between data points is calculated and produces clusters with different shapes and separations between each other. As it has been raised in Section “Background and related work”, despite the widespread use of Euclidean distance in indexing structures and algorithms, relying on it as the default option when using the $k$-means algorithm can lead to poor results.
Previous research [22] suggests that inherent characteristics of high-dimensional spaces, such as high dispersion, outliers or concentration of points around specific values, hinder the discriminatory power of the $\ell_{p}norm$ for higher values of $p$. This situation would provide undue importance to the distant (sparse/noisy) dimensions and poorly discriminate between further and nearer neighbours. Therefore, for a given high-dimensional problem, it may be preferable to use Minkowski distances with lower values of $p$, so the $\ell_{1}$ distance metric (Manhattan distance) would be the best choice for high-dimensional applications, followed by the Euclidean distance ($\ell_{2}$), then the $\ell_{3}$ metric, up to the Chebyshev distance ($\ell_{\infty}$).
Previous algorithms for approximate similarity search, such as MASK [2], have relied on the Euclidean distance to calculate relative dissimilarities between points. This strategy follows the conventional inclination to associate distance between points with the Euclidean metric, which aligns with the distance concept commonly applied in geospatial problems. However, the potential for achieving improved results, coupled with the lack of prior empirical studies that validate this proposal and the need for learning paradigms applicable to a broader range of dissimilarity measures [5], justify our efforts to explore alternative Minkowski distances for data indexing. This endeavour aims to foster a deeper understanding of the impact of using alternative distance metrics with high-dimensional data.
Nevertheless, while it may seem theoretically plausible to employ $k$-means or its enhanced version, $k$-means++, with several distance metrics, the practical implementation reveals that the algorithm’s internal design is intertwined with the Euclidean distance. This alignment relies on the algorithm’s iterative optimisation process when it searches for centroids that minimise the total squared distances of points within their respective clusters[3]. The sum of the Squared Error (SSE), calculated as the sum of squared Euclidean distances between each data point and the centroid of its assigned cluster, underscores the algorithm’s inherent reliance on Euclidean geometry for clustering optimisation [28].
As we have already demonstrated, assuming that widely adopted mechanisms are the most suitable or the only viable option can lead to self-imposed limitations that generate sub-optimal outcomes. Instead, our objective is to steer clear of the misconception that certain implementations are the best or the only conceivable choice simply because they are the most common choices.
At this juncture, GMASK can integrate an alternative clustering algorithm that leverages the opportunities provided by diverse distance metrics for data indexing. The $k$-means algorithm defines clusters based on centroids minimising the sum of squared errors in distances from all other points in the same cluster. In contrast, alternative methods like $k$-medoids calculate clusters around the most representative data element, which is determined by minimising the sum of pairwise dissimilarities among data points the same cluster.
3.5 k-medoids as a generalisation of k-means
Several generalisations of the $k$-means algorithm have been proposed to broaden the scope of applicable data types [29], reflecting the diversity found in nature prototype points, such as subsets of objects, hyperplanes, etc.
The first paper proposes a general criterion and its generalised $k$-means method is [30]. This approach starts from a $n$ x $n$ dissimilarity matrix $(d_{\ell}t)$ for $n$ factories $O=1,...,n$ in an industrial network where $d_{\ell}t$ are the minimum road transportation costs between $\ell$ and $t$. The goal is to partition the set of factories into $k$ classes $(C_{1},...,C_{k})$ and to find a selection $Z=(z_{1},...,z_{k})$ of $k$ factories as ’supply points’ such that, when supplying all factories of the class $C_{i}$ from the supply point $z_{i}\in O$, the overall transport costs are minimised, where $d(\ell,z_{i})=d_{\ell},z_{i}$ means the dissimilarity between the factory (object) $\ell$ and the factory (supply point) $z_{i}\in O$. So the family $P$ of admissible ’prototypes’ consists of all singletons from $O$ and the class prototypes, meaning ’most cheapest supply point’ in $C_{i}$, are determined by:
$Q(C_{i},z):=\sum_{\ell\in C_{i}}d(\ell,z)\to\min_{z\in C_{i}}$ (3)
Many authors, including, among others, Diday [31, 32, 33], Bock [34] or Govaert [35], have followed the generalised clustering approach in various settings and numerous variations, obtaining a plethora of generalised $k$-means algorithms.
Kaufmann and Rousseeuw [36] first propose the popular Partitioning Around Medoids (PAM) algorithm, also referred to as $k$-medoids. In $k$-medoids, data is modelled by using $k$ representative objects $m_{i}$ called medoids, that serve as prototypes for the clusters (formerly called classes) which are obtained by using the absolute error criterion (’total deviation’, TD). Since the TD is defined as the sum of dissimilarities of each point $x_{j}\in C_{i}$ to the medoid $m_{i}$ of its cluster, the medoid of a cluster $C_{i}$ is in fact the object with the smallest sum of dissimilarities (or equivalently, the smallest average) to all other objects in the cluster.
$medoid(C_{i}):=argmin_{x_{i}\in C}\sum_{{x_{j}\in C}}d(x_{i},x_{j})$ (4)
This method is suitable for non-Euclidean distances and a wide range of data since it does not force the dissimilarity measure to adhere strictly to metric properties, and only requires a proximity measure between any pair of objects. In addition, while $k$-means defines a prototype in terms of a centroid, which is usually the mean of a group of points that may not correspond to an element in the dataset, $k$-medoids defines a prototype in terms of a medoid, an actual data point identified as the most representative one for a group of points.
Regarding $k$-medoid implementations, Program PAM [36] consists of two algorithms, BUILD to choose an initial clustering, and SWAP to improve the clustering towards a local optimum, since finding the global optimum of the $k$-medoids problem is, unfortunately, NP-hard. The algorithms require a dissimilarity matrix, which requires $O(n^{2})$ memory and at least $O(n^{2}d)$ time to be computed. On top of this, the process of updating medoids involves evaluating the cost of swapping non-medoid points to ensure that the sum of pairwise dissimilarities within clusters (rather than the sum of squared Euclidean distances) is minimised, making this method more robust to noise and outliers compared to the alternatives based on $k$-means. Recently, FasterPAM [37] introduced a speedup for larger $k$, achieving an $O(k)$-fold acceleration in the second SWAP phase of the algorithm, but still providing the same results as the original PAM algorithm.
Originally FasterPAM was implemented in Java and published within the open-source library ELKI [38]. However, to integrate it with GMASK implementation, developed entirely in Python, we will employ the fast $k$-medoids Python implementation [39] version provided by the $kmedoids$ package  [^1] https://pypi.org/project/kmedoids. FasterPAM let us upgrade our method by making it compatible with any distance function and a wider range of problems.
Experimental results verify the applicability of GMASK with real datasets, improving the performance of alternative algorithms and providing better results for approximate similarity search queries.
3.6 Performance evaluation
The main advantage of GMASK is providing a fast response to queries at the expense of returning approximate results. Thus, a straightforward strategy to assess the performance of this algorithm is to undertake an exhaustive search of all elements in a given dataset $X$. Then, the proportion of data points that were correctly retrieved can be determined. This is the common approach to all evaluation experiments developed in this study.
A comprehensive series of experiments are conducted across datasets of diverse types, sizes, dimensionality, and sparsity to evaluate the performance of GMASK, which introduces $k$-medoids for the bottom-top construction of the index tree. This novelty enables its parametrisation with several $\ell_{p}$ norms as outlined in “k-medoids as a generalisation of k-means”. These results will be compared with those obtained using other state-of-the-art algorithms. As mentioned above, our objective is to understand better the impact of utilising distances other than the Euclidean metric. Through these experiments, we aim to demonstrate that data indexing using non-Euclidean distances is not only feasible but also has the potential to yield improved outcomes.
3.7 Benchmark algorithms for similarity search
To evaluate our thesis, we compare GMASK performance in approximate nearest neighbour search with respect to alternative MAM.
3.7.1 KD-Tree
The KD-Tree (K-Dimensional Tree) [40] is an exact similarity search method used to find the nearest neighbours to a given query point. It constructs a binary tree structure by recursively dividing the data based on the maximum spread along different dimensions and the search algorithm efficiently narrows down the search space by traversing the tree. In low-dimensional spaces, KD-trees can achieve sublinear search times. However, as the dimensionality increases, the performance of the KD-tree can degrade due to the Curse of Dimensionality phenomenon.
As it performs an exact search, the solutions provided by this method are used as a baseline to evaluate the performance of other approximate algorithms. We shall compare the k-nearest neighbours obtained by each method against those obtained using the KD-Tree to assess their accuracy.
3.7.2 FLANN
FLANN (Fast Library for Approximate Nearest Neighbours) [11] is an efficient method widely employed for approximate similarity search in large-scale and high-dimensional datasets, extensively used in machine learning, computer vision, and data mining. FLANN employs techniques such as random projection and hierarchical subdivision to construct an index structure that enables faster search operations. It offers several index types, including KD-tree, $k$-means tree, and hierarchical clustering tree, which are suitable for diverse data types and query requirements. While FLANN has implementations in several programming languages, we use the Python bindings provided by the pyflann  [^2] https://github.com/primetang/pyflann library.
3.7.3 PyNNDescent
The PyNNDescent algorithm is an approximate similarity search method [41] implemented in Python  [^3] https://pynndescent.readthedocs.io/en/latest/. It is based on the NNDescent algorithm and used for efficient nearest neighbour search in large-scale datasets. PyNNDescent constructs a graph-based index by connecting each data point to its approximate nearest neighbours. It iteratively improves the graph structure through neighbour descent steps, employing techniques such as random initialisation, multi-tree searching, and pruning. The algorithm also provides parameters to control the trade-off between search accuracy and computational cost. PyNNDescent is particularly useful for high-dimensional datasets and is widely used in machine learning, data mining, and information retrieval tasks.
3.8 Datasets
Several experiments have been conducted to evaluate the performance of GMASK, using four datasets diverse in nature and public access. Table 1 briefly summarises their main features.
3.8.1 Municipalities
The Municipalities dataset is a two-dimensional geospatial dataset that contains the coordinates (latitude and longitude) of each of the 8,130 municipalities that make up the Spanish territory, as depicted in Figure 3. It has been extracted from one of the tables provided in the ’2020 Resumen Descriptivo del Nomenclátor de Municipios y Entidades de Población’ (Descriptive Summary of the Nomenclature of Municipalities and Population Entities), called MUNICIPIOS [^4] https://ign.es/web/rcc-nomenclator-nacional. This report is compiled annually by the Instituto Geográfico Nacional (National Geographic Institute), a public institution under the Spanish Ministry of Transport and Sustainable Mobility. It includes updated geographic and administrative data on all Spanish municipalities, such as their name, province, geographic location coordinates, perimeter or population, among others.
Figure 3: Visualisation for dataset Municipalities
Although this is a two-dimensional dataset with a low dispersion, it is an exceptionally intriguing case due to its intuitive representation and the presence of outliers, such as the Canary Islands and the Balearic Islands. The location of these islands deviates from the rest of the points significantly and poses a challenge to any similarity search method applied to this dataset.
3.8.2 MNIST
The MNIST_784 dataset  [^5] http://yann.lecun.com/exdb/mnist/ is a widely used machine learning benchmark dataset whose simplicity makes it an ideal starting point for algorithm development and evaluation. Firstly, it consists of 60,000 training images and 10,000 test images, a relatively small size that makes possible to carry out computationally efficient experimentation and prototyping. Secondly, it comprises grayscale images of handwritten digits ranging from 0 to 9, making it visually straightforward and easy to work with. Moreover, each image is 28x28 pixels in size, resulting in a total of 784 dimensions when flattened, allowing researchers to explore algorithms in the context of high-dimensional data. Many well-established techniques and algorithms have been evaluated on MNIST [42, 43], making it a reference point for evaluating the performance of approximate similarity search methods [16, 44].
3.8.3 GLOVE
The third experiment involves the GLOVE_100 dataset [^6] https://nlp.stanford.edu/data/glove.6B.zip, which contains one million pre-trained Global Vectors for Word Representation (GloVe) embeddings that capture semantic relationships between words. These 100-dimensional embedding vectors were pre-trained on the combined Wikipedia 2014 + Gigaword 5th Edition corpora (6B tokens, 400K vocabulary) and cover a wide range of English words, enabling comprehensive assessments of approximate similarity search methods across diverse vocabulary and word relationships. Given their widespread adoption as a benchmark in the natural language processing community, the use of GloVe embeddings as a standardised dataset provides valuable insights into the method effectiveness in retrieving similar items based on their semantic similarity [16].
3.8.4 NYtimes
The NYtimes dataset, part of the Bags of Words dataset from the UCI repository [^7] https://archive.ics.uci.edu/dataset/164/bag+of+words, comprises a collection of New York Times news articles represented as a bag of words. Each document in the dataset is associated with a set of word occurrences, where the dimensions represent unique words extracted from the articles. The dataset is organised as a document–word matrix, where each row corresponds to a document and each column corresponds to a word. The values in the matrix indicate the frequency of each word occurring in the respective document. Preprocessing steps include tokenization, removal of stopwords, and vocabulary truncation, with only words occurring more than ten times retained.
The reduced version of the NYtimes employed on this work was generated by building tf-idf descriptors from the bag-of-words version, and embedding them into a lower dimensional space using the Johnson-Lindenstrauss Transform [45].
Dataset Label n Dimensionality High Sparsity Data Type
Municipalities $\mathcal{X}_{muni}$ 8,130 2 No Geospatial
MNIST $\mathcal{X}_{MNIST}$ 69,000 784 Yes Image
GLOVE $\mathcal{X}_{GLOVE}$ 1,000,000 100 No Text
NYtimes $\mathcal{X}_{NYtimes}$ 290,000 256 No Text
Table 1: Descriptive traits characterising the datasets included in experiments.
4 Experimental results
In order to show the versatility of GMASK, we present the analysis of four different datasets: municipalities, MNIST, GLOVE and NYtimes. This study consists of a neighbours search applying Euclidean distance, Manhattan distance and Chebyshev distance. The main objective is to measure the accuracy of the GMASK algorithm with different distances and to observe if any distance is more adequate for some dataset, depending on its characteristics. For each dataset, we compare the results obtained by GMASK with two other algorithms: PYNN and FLANN. Both PYNN and GMASK are able to work with these three distances, however, FLANN does not work with Chebyshev distance. Since MASK is a specific case of GMASK only able to work with Euclidean distance, its inclusion in these experiments has not been considered.
Our approach to evaluating GMASK performance is conceived to closely emulate real-world conditions. The datasets used in this work are not artificially generated, and no preprocessing has been applied. The transformation and cleaning of the collected data is limited to dividing the original dataset into two distinct subgroups, the training and testing datasets. We also check that identical elements do not appear in both subgroups, thereby intensifying the complexity in the search for $k$ approximate neighbours.
All experiments have been performed on a server equipped with 2 AMD EPYC 7451 microprocessors (24 cores/48 threads, 2.30 GHz, 64 MB L3 cache), 128 GB of DDR4 RAM and an Intel D3-S4510 SSD (capacity 480 GB) for secondary storage. The experimental results are presented in the following sections.
4.1 Municipalities
The municipalities dataset is an interesting case due to the geographical distribution of points. In Figure 3, we can identify three main groups of points: the largest region corresponds to the peninsular territory of Spain, whereas the group of points located at the right of the map corresponds to the Balearic Islands and the group of points located at the bottom left corner of the map corresponds to the Canary Islands. The last two groups of points increase the difficulty of the analysis because they are far from the rest of the points, so we can consider them as outliers. Despite this dataset containing additional information, we only consider two variables: latitude and longitude.
Figure 4 shows the recall comparison between FLANN, PYNN and GMASK. We use a heatmap to represent the recall value obtained for each algorithm with each distance in the search of 5, 10 and 15 neighbours. The recall obtained in this case for each method is very similar, and all of them depict good recall values, over 84%. It is noticeable that FLANN recall decreases by about 10% with the number of neighbours. Regardless of the distance used and the number of neighbours intended to search, GMASK is the only method that provides a 100% recall in all tests.
Figure 4: Heatmap comparing the Recall (percentage of true nearest neighbours found) obtained by each algorithm and distance in approximate nearest neighbour search, for different number of neighbours, with dataset $\mathcal{X}_{muni}$.
4.2 MNIST
Given its high dimensionality and sparsity, MNIST has become a widespread benchmark in similarity search research. The following experiment evaluates the recall obtained in searching neighbours in this dataset, applying Manhattan, Euclidean and Chebyshev distance with FLANN, PYNN and GMASK methods. Figure 5 presents the results of this analysis.
From this comparison, we can deduce that the results are highly comparable in the recall when applying PYNN and GMASK methods with Euclidean and Manhattan distances, respectively. Therefore, the combination of both distances and these methods suits this case study. However, this is not the case for Chebyshev distance since the recall is lower than 70%. There is also a meaningful reduction in FLANN recall when the number of neighbours increases. Regardless of the distance, GMASK yields a higher average recall than other methods.
Figure 5: Heatmap comparing the Recall (percentage of true nearest neighbours found) obtained by each algorithm and distance in approximate nearest neighbour search, for different number of neighbours, with dataset $\mathcal{X}_{MNIST}$.
4.3 GLOVE
Hereafter, the third experiment involves the analysis of the GLOVE dataset, a high dimensional datasets that contains the largest number of elements among the datasets included in these experiments.
Figure 6 depicts the recall results obtained for each method and distance. Once again, the luck of efficiency is evident in FLANN´s results. The best recall obtained with FLANN is 12% when searching for 5 neighbours applying Euclidean distance. PYNN method provides good accuracy when Manhattan distance is used. If we focus on PYNN method and compare the recall for the three distances, we can see that Manhattan distance provides the best results and Chebyshev distance provides the worst recalls. However, in the case of GMASK, the recall is 100% per each distance and with independence of the number of neighbours searched.
Figure 6: Heatmap comparing the Recall (percentage of true nearest neighbours found) obtained by each algorithm and distance in approximate nearest neighbour search, for different number of neighbours, with dataset $\mathcal{X}_{GLOVE}$.
4.4 NYtimes
Finally, the last experiment replicates the same analysis using a dataset characterised by high dimensionality, featuring a data-to-dimensionality ratio lower than that of GLOVE but higher than MNIST: the NYtimes dataset.
Figure 7 illustrates the recall accomplished by FLANN, PYNN and GMASK. There is a noticeable difference between GMASK and the other methods, since the recall values for FLANN and PYNN are below 50%. However, GMASK consistently outperforms the other methods regardless of the distance chosen. Once again, FLANN exhibits the poorest performance.
Therefore, while the most remarkable insight obtained from this experiment is the superior performance of the Chebyshev distance metric compared to others, it is crucial to highlight that, in this case, GMASK nearly doubles the recall provided by PYNN, achieving an 82% recall when searching for the 5 nearest neighbours.
Through this experiment, we empirically demonstrate what was previously stated theoretically in earlier sections of this manuscript: adopting the Euclidean distance as the default option in indexing structures and algorithms is a self-imposed restriction that can lead to sub-optimal results.
Figure 7: Heatmap comparing the Recall (percentage of true nearest neighbours found) obtained by each algorithm and distance in approximate nearest neighbour search, for different number of neighbours, with dataset $\mathcal{X}_{NYtimes}$.
5 Discussion
Given the notable prevalence of Euclidean distance in similarity searches in the previous literature, as well as the limitations detected in the widely adopted $k$-means algorithm, which is strongly linked to this metric, there is a growing need to explore the use of alternative metrics when addressing these types of problems. Stemming from this need, GMASK makes the accommodation of arbitrary distances feasible by integrating the $k$-medoids algorithm as a clustering algorithm, allowing us to leverage the advantages derived from using several distance metrics on high-dimensional datasets.
It is crucial to emphasise that the algorithm’s generalisation, achieved through the integration of $k$-medoids, significantly broadens the range of applicable distances while preserving the structural integrity and properties of the original algorithm. Thus, GMASK maintains crucial characteristics, including the desirable scalability property facilitated by the bottom-top construction of the search tree, retaining a competitive edge over other state-of-the-art algorithms that require centralising all data on a single server.
As elaborated in Section “Experimental results”, our findings demonstrate the efficiency of this new approach, thereby supporting the hypothesis that alternative distances beyond the Euclidean metric may yield superior results in solving nearest-neighbour queries, contingent upon the specific characteristics of the dataset. Therefore, it becomes evident that the outcomes achieved using the proposed method surpass those obtained through two other widely adopted state-of-the-art algorithms, which are distinct from each other and significantly differ from GMASK.
These findings would also be in line with previous publications [22] suggesting that, for a given high-dimensional problem, using an $\ell_{p}$ norm with lower values of $p$ may be advantageous in assessing dissimilarity between query points. Thus, while Aggarwal carries out a theoretical approach using synthetic data, our research delves into this hypothesis by generalising an algorithm that has been tested over a range of real-world datasets of diverse nature. However, further empirical research is required to validate this observed behaviour more precisely.
GMASK provides an avenue for future research to integrate distance functions beyond Minkowski, thus providing many opportunities to extend this research line. For instance, it would be interesting to explore the suitability of distance functions for specific purposes, such as cosine similarity for textual data or the Jaccard distance for categorical data, to address particular problems.
Furthermore, although GMASK has yielded notably satisfactory experimental results, there still exists potential for enhancing this implementation to achieve further optimisation in computational methods, thereby increasing overall performance. To this end, the proposal suggests exploring the incorporation of a just-in-time (JIT) compiler for Python, such as Numba  [^8] https://numba.pydata.org—employed by other algorithms such as PYNNDescent—, or considering the adaptation and port of GMASK to a language renowned for its strong performance in building scalable and concurrent systems like Go  [^9] https://go.dev.
6 Conclusions and future work
Approximate similarity search is essential for many application areas where large and heterogeneous datasets arise, often exhibiting high dimensionality and sparsity. In many cases, these datasets must be stored in distributed systems to become manageable, imposing additional restrictions that rule out algorithms conceived for cases where data must reside on a single machine.
This paper presents the GMASK algorithm for approximate similarity search in distributed systems that can accept arbitrary distance functions. One of the main novelties is adopting the $k$-medoids (PAM) algorithm, which is not bounded to the Euclidean distance, to achieve efficient solutions. Furthermore, using FastPAM, a highly efficient implementation of the PAM algorithm to solve the $k$-medoids problem, GMASK can improve its efficiency with respect to other previous algorithms.
Experiments conducted on real-world datasets with different properties confirm that GMASK, based on $k$-medoids, performs better in solving nearest-neighbour queries than other frequently used algorithms, using other Minkowski dissimilarity functions different from the Euclidean distance. These results illustrate that indexing data with non-Euclidean distances is not only viable but also holds the potential to enhance search outcomes substantially. Overlooking alternative distances would represent a voluntary limitation that may hinder us from achieving the optimal results.
GMASK lays out new interesting routes for further research. Once the requirement of using the Euclidean distance has been lifted, additional studies should be conducted to evaluate GMASK performance in other settings where alternative distances can be applied, such as pseudo-metrics (Minkowski distances with $p<1$) or non-metric dissimilarities. Besides, using arbitrary distances will enable us to consider datasets with qualitative features exclusively or a combination of quantitative and qualitative features. Many state-of-the-art algorithms and standard dissimilarity measures cannot be directly applied in these cases since they are incompatible with categorical features.